# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v2

@@ D.3 Effect of CoT Reasoning
GPT-4 Omni GPT-4o-Mini $55.77$ $81.47$ $-16.34$ $-10.71$ $-19.73$ $-8.84$ $-13.89$ $-8.91$ $-12.98$ $-2.19$ $-15.62$ $-7.75$
GPT-4o $68.78$ $91.79$ $-13.55$ $-4.53$ $-13.27$ $-4.61$ $-13.55$ $-5.16$ $-9.28$ $-1.25$ $-12.45$ $-3.87$
Claude Series Claude-3.5-Haiku-20241022 $57.73$ $72.95$ $-16.68$ $-15.56$ $-7.14$ $-7.74$ $-5.79$ $-6.33$ $-21.50$ $-10.40$ $-15.33$ $-10.43$
Claude-3.5-Sonnet-20241022 $72.35$ $86.00$ $-16.03$ $-6.49$ $-6.28$ $-4.53$ $-6.80$ $-5.47$ $-18.45$ $-6.80$ $-14.26$ $-5.90$
Gemini Series Gemini-1.5-Flash $51.73$ $75.22$ $-8.57$ $-13.76$ $-10.53$ $-16.34$ $-16.71$ $-31.74$ $-6.25$ $-1.56$ $-9.26$ $-15.34$
Gemini-2.0-Flash $65.39$ $89.13$ $-23.92$ $-5.55$ $-23.72$ $-5.63$ $-25.91$ $-12.74$ $-8.18$ $-1.80$ $-20.21$ $-6.25$
Gemini-1.5-Pro-002 $63.17$ $87.18$ $-13.27$ $-6.49$ $-16.63$ $-10.40$ $-16.05$ $-12.82$ $-11.05$ $-3.36$ $-13.48$ $-8.01$
DeepSeek DeepSeek-V3 $67.84$ $89.52$ $-14.52$ $-8.91$ $-17.54$ $-15.87$ $-12.98$ $-16.81$ $-12.09$ $-3.13$ $-14.10$ $-10.81$
LLaMA Series LLaMA-3.1-8B-Instruct $35.47$ $44.72$ $-7.90$ $-12.35$ $-8.76$ $-9.38$ $-7.01$ $-9.62$ $-8.84$ $-3.91$ $-8.13$ $-8.96$
LLaMA-3.1-70B-Instruct $51.89$ $69.43$ $-9.77$ $-12.43$ $-11.44$ $-16.11$ $-15.87$ $-21.42$ $-8.37$ $-4.69$ $-10.39$ $-13.35$
LLaMA-3.1-405B-Instruct $58.72$ $78.42$ $-11.02$ $-10.63$ $-5.39$ $-8.05$ $-8.26$ $-12.20$ $-7.61$ $-5.86$ $-9.12$ $-9.21$
LLaMA-3.3-70B-Instruct $55.64$ $76.86$ $-10.14$ $-8.37$ $-8.26$ $-5.94$ $-9.67$ $-8.52$ $-6.36$ $-3.67$ $-8.90$ $-6.68$
Qwen Series Qwen2.5-7B-Instruct $42.59$ $58.17$ $-15.06$ $-12.67$ $-21.66$ $-12.74$ $-15.32$ $-16.26$ $-15.92$ $-4.07$ $-16.14$ $-11.34$
Qwen2.5-14B-Instruct $48.42$ $70.68$ $-17.46$ $-16.34$ $-18.09$ $-24.94$ $-16.00$ $-30.65$ $-6.91$ $-7.66$ $-14.72$ $-19.33$
Qwen2.5-32B-Instruct $59.89$ $79.44$ $-15.22$ $-13.84$ $-14.07$ $-14.78$ $-14.15$ $-19.23$ $-11.81$ $-4.22$ $-14.09$ $-12.86$
Qwen2.5-72B-Instruct $58.14$ $82.80$ $-13.94$ $-10.48$ $-8.99$ $-5.63$ $-14.44$ $-13.45$ $-14.36$ $-5.47$ $-13.49$ $-8.80$
Qwen2.5-Coder-32B-Instruct $63.10$ $84.75$ $-15.27$ $-12.43$ $-17.49$ $-16.11$ $-13.37$ $-16.89$ $-8.99$ $-7.90$ $-13.74$ $-13.13$
Average $\mathbf{-14.04}$ $\mathbf{-10.68}$ $\mathbf{-13.47}$ $\mathbf{-11.04}$ $\mathbf{-13.28}$ $\mathbf{-14.60}$ $\mathbf{-11.11}$ $\mathbf{-4.58}$ $\mathbf{-13.14}$ $\mathbf{-10.12}$
Table 15: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Absolute Version]
Model Series Model Name VAN ALL MDC MPS MHC
@@
GPT-4 Omni GPT-4o-Mini $79.88$ $84.13$ $-10.80\%$ $-16.87\%$ $-10.80\%$ $-10.92\%$ $-13.15\%$ $-7.44\%$ $-3.29\%$ $-1.74\%$
GPT-4o $90.25$ $94.36$ $-2.77\%$ $-8.41\%$ $-3.60\%$ $-7.30\%$ $-4.71\%$ $-7.08\%$ $-2.49\%$ $+0.44\%$
Claude Series Claude-3.5-Haiku-20241022 $71.62$ $75.16$ $-16.06\%$ $-29.72\%$ $-9.95\%$ $-11.67\%$ $-9.77\%$ $-6.94\%$ $-11.69\%$ $-18.33\%$
Claude-3.5-Sonnet-20241022 $83.75$ $89.77$ $-4.33\%$ $-12.56\%$ $-5.97\%$ $-4.19\%$ $-4.48\%$ $-9.30\%$ $-6.12\%$ $-10.70\%$
Gemini Series Gemini-1.5-Flash $74.62$ $76.20$ $-17.09\%$ $-20.27\%$ $-13.23\%$ $-35.62\%$ $-34.00\%$ $-55.62\%$ $-2.01\%$ $-2.19\%$
Gemini-2.0-Flash $87.25$ $92.28$ $-5.01\%$ $-8.14\%$ $-6.02\%$ $-6.79\%$ $-7.59\%$ $-24.89\%$ $-2.29\%$ $-1.58\%$
Gemini-1.5-Pro-002 $84.50$ $91.65$ $-6.51\%$ $-8.88\%$ $-10.80\%$ $-13.67\%$ $-11.98\%$ $-18.91\%$ $-4.44\%$ $-2.96\%$
DeepSeek DeepSeek-V3 $88.00$ $92.07$ $-7.39\%$ $-14.06\%$ $-13.92\%$ $-23.81\%$ $-15.77\%$ $-23.58\%$ $-4.26\%$ $-2.27\%$
LLaMA Series LLaMA-3.1-8B-Instruct $46.75$ $41.34$ $-28.34\%$ $-26.26\%$ $-19.52\%$ $-23.74\%$ $-22.19\%$ $-20.20\%$ $-5.61\%$ $-14.65\%$
LLaMA-3.1-70B-Instruct $70.75$ $67.22$ $-15.19\%$ $-22.67\%$ $-18.90\%$ $-30.75\%$ $-24.56\%$ $-41.93\%$ $-5.65\%$ $-8.70\%$
LLaMA-3.1-405B-Instruct $78.88$ $77.66$ $-9.83\%$ $-19.89\%$ $-7.29\%$ $-15.32\%$ $-10.14\%$ $-24.73\%$ $-3.96\%$ $-13.44\%$
LLaMA-3.3-70B-Instruct $76.12$ $78.08$ $-8.21\%$ $-15.24\%$ $-5.42\%$ $-11.50\%$ $-5.75\%$ $-19.79\%$ $-1.81\%$ $-9.63\%$
Qwen Series Qwen2.5-7B-Instruct $57.25$ $59.71$ $-17.47\%$ $-28.67\%$ $-23.80\%$ $-18.88\%$ $-30.35\%$ $-24.13\%$ $-7.42\%$ $-6.29\%$
Qwen2.5-14B-Instruct $67.25$ $76.41$ $-20.26\%$ $-27.32\%$ $-36.06\%$ $-34.15\%$ $-43.49\%$ $-43.17\%$ $-11.71\%$ $-9.56\%$
Qwen2.5-32B-Instruct $77.62$ $82.46$ $-14.01\%$ $-22.78\%$ $-16.26\%$ $-22.28\%$ $-22.38\%$ $-27.09\%$ $-6.60\%$ $-3.29\%$
Qwen2.5-72B-Instruct $78.25$ $90.40$ $-8.31\%$ $-18.94\%$ $-7.67\%$ $-5.54\%$ $-16.61\%$ $-15.70\%$ $-5.27\%$ $-8.55\%$
Qwen2.5-Coder-32B-Instruct $84.12$ $85.80$ $-10.10\%$ $-22.14\%$ $-16.79\%$ $-22.63\%$ $-13.97\%$ $-29.68\%$ $-8.02\%$ $-11.44\%$
Average $-11.86\%$ $-18.99\%$ $-13.29\%$ $-17.57\%$ $-17.11\%$ $-23.54\%$ $-5.45\%$ $-7.35\%$
$\mathbf{-14.53}\%$ $\mathbf{-14.90}\%$ $\mathbf{-19.52}\%$ $\mathbf{-6.16}\%$
Table 16: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Relative Version]
Model Series Model Name VAN ALL MDC MPS MHC
@@
GPT-4 Omni GPT-4o-Mini $79.88$ $84.13$ $-8.62$ $-14.20$ $-8.62$ $-9.19$ $-10.50$ $-6.26$ $-2.62$ $-1.46$
GPT-4o $90.25$ $94.36$ $-2.50$ $-7.93$ $-3.25$ $-6.89$ $-4.25$ $-6.68$ $-2.25$ $+0.42$
Claude Series Claude-3.5-Haiku-20241022 $71.62$ $75.16$ $-11.50$ $-22.34$ $-7.12$ $-8.77$ $-7.00$ $-5.22$ $-8.38$ $-13.78$
Claude-3.5-Sonnet-20241022 $83.75$ $89.77$ $-3.62$ $-11.27$ $-5.00$ $-3.76$ $-3.75$ $-8.35$ $-5.12$ $-9.60$
Gemini Series Gemini-1.5-Flash $74.62$ $76.20$ $-12.75$ $-15.45$ $-9.88$ $-27.14$ $-25.38$ $-42.38$ $-1.50$ $-1.67$
Gemini-2.0-Flash $87.25$ $92.28$ $-4.38$ $-7.52$ $-5.25$ $-6.26$ $-6.62$ $-22.96$ $-2.00$ $-1.46$
Gemini-1.5-Pro-002 $84.50$ $91.65$ $-5.50$ $-8.14$ $-9.12$ $-12.53$ $-10.12$ $-17.33$ $-3.75$ $-2.71$
DeepSeek DeepSeek-V3 $88.00$ $92.07$ $-6.50$ $-12.94$ $-12.25$ $-21.92$ $-13.88$ $-21.71$ $-3.75$ $-2.09$
LLaMA Series LLaMA-3.1-8B-Instruct $46.75$ $41.34$ $-13.25$ $-10.86$ $-9.12$ $-9.81$ $-10.38$ $-8.35$ $-2.62$ $-6.05$
LLaMA-3.1-70B-Instruct $70.75$ $67.22$ $-10.75$ $-15.24$ $-13.38$ $-20.67$ $-17.38$ $-28.18$ $-4.00$ $-5.85$
LLaMA-3.1-405B-Instruct $78.88$ $77.66$ $-7.75$ $-15.45$ $-5.75$ $-11.90$ $-8.00$ $-19.21$ $-3.12$ $-10.44$
LLaMA-3.3-70B-Instruct $76.12$ $78.08$ $-6.25$ $-11.90$ $-4.12$ $-8.98$ $-4.38$ $-15.45$ $-1.38$ $-7.52$
Qwen Series Qwen2.5-7B-Instruct $57.25$ $59.71$ $-10.00$ $-17.12$ $-13.62$ $-11.27$ $-17.38$ $-14.41$ $-4.25$ $-3.76$
Qwen2.5-14B-Instruct $67.25$ $76.41$ $-13.62$ $-20.88$ $-24.25$ $-26.10$ $-29.25$ $-32.99$ $-7.88$ $-7.31$
Qwen2.5-32B-Instruct $77.62$ $82.46$ $-10.88$ $-18.79$ $-12.62$ $-18.37$ $-17.38$ $-22.34$ $-5.12$ $-2.71$
Qwen2.5-72B-Instruct $78.25$ $90.40$ $-6.50$ $-17.12$ $-6.00$ $-5.01$ $-13.00$ $-14.20$ $-4.12$ $-7.72$
Qwen2.5-Coder-32B-Instruct $84.12$ $85.80$ $-8.50$ $-19.00$ $-14.12$ $-19.42$ $-11.75$ $-25.47$ $-6.75$ $-9.81$
Average $-8.40$ $-14.48$ $-9.62$ $-13.41$ $-12.38$ $-18.32$ $-4.04$ $-5.50$
$\mathbf{-10.68}$ $\mathbf{-11.04}$ $\mathbf{-14.60}$ $\mathbf{-4.58}$
D.4 Impact on Input Prediction
In §4.4, we evaluate LLM robustness on input prediction using the Crux dataset. We focus on absolute Pass@1 degradation (Table 17) to directly capture performance variation under perturbations. For completeness, we also report relative degradation in Table 5.
Table 17: Comparison of absolute Pass@1 degradation under Crux for input prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Relative Version]
Model Series Model Name VAN ALL MDC MPS MHC

### v3

@@ E.2 Additional Performance Degradation Tables for Output Prediction (CoT Inference)
GPT-4 Omni GPT-4o-Mini $55.8$ $81.5$ 29.4-29.4 13.1-13.1 35.5-35.5 10.8-10.8 25.0-25.0 11.0-11.0 23.8-23.8 2.7-2.7 28.4-28.4 9.4-9.4
GPT-4o $68.8$ $91.8$ 20.0-20.0 4.9-4.9 19.6-19.6 5.0-5.0 19.9-19.9 5.6-5.6 14.0-14.0 1.4-1.4 18.4-18.4 4.2-4.2
Claude Series Claude-3.5-Haiku-20241022 $57.7$ $72.9$ 28.9-28.9 21.2-21.2 12.4-12.4 10.6-10.6 10.0-10.0 8.7-8.7 37.2-37.2 14.2-14.2 22.1-22.1 13.7-13.7
Claude-3.5-Sonnet-20241022 $72.3$ $86.0$ 22.0-22.0 7.4-7.4 8.7-8.7 5.3-5.3 9.4-9.4 6.3-6.3 25.3-25.3 7.8-7.8 16.3-16.3 6.7-6.7
Gemini Series Gemini-1.5-Flash $51.7$ $75.2$ 16.3-16.3 18.3-18.3 21.3-21.3 21.6-21.6 32.9-32.9 42.1-42.1 12.9-12.9 2.1-2.1 20.8-20.8 21.0-21.0
Gemini-2.0-Flash $65.4$ $89.1$ 36.6-36.6 6.2-6.2 36.2-36.2 6.3-6.3 39.6-39.6 14.1-14.1 12.5-12.5 2.0-2.0 31.2-31.2 7.1-7.1
Gemini-1.5-Pro-002 $63.2$ $87.2$ 21.1-21.1 7.4-7.4 26.7-26.7 11.9-11.9 25.9-25.9 14.6-14.6 18.4-18.4 3.9-3.9 23.0-23.0 9.4-9.4
DeepSeek DeepSeek-V3 $67.8$ $89.5$ 21.4-21.4 9.9-9.9 25.9-25.9 17.6-17.6 19.1-19.1 18.7-18.7 17.8-17.8 3.5-3.5 21.1-21.1 12.4-12.4
LLaMA Series LLaMA-3.1-8B-Instruct $35.5$ $44.7$ 22.2-22.2 27.6-27.6 24.7-24.7 21.1-21.1 19.8-19.8 21.4-21.4 25.1-25.1 9.0-9.0 23.0-23.0 19.8-19.8
LLaMA-3.1-70B-Instruct $51.9$ $69.4$ 18.7-18.7 18.0-18.0 22.5-22.5 23.3-23.3 31.2-31.2 31.1-31.1 17.3-17.3 6.8-6.8 22.4-22.4 19.8-19.8
LLaMA-3.1-405B-Instruct $58.7$ $78.4$ 18.6-18.6 13.6-13.6 9.6-9.6 10.3-10.3 14.5-14.5 15.6-15.6 13.9-13.9 7.5-7.5 14.2-14.2 11.8-11.8
LLaMA-3.3-70B-Instruct $55.6$ $76.9$ 18.4-18.4 10.8-10.8 15.2-15.2 7.7-7.7 17.9-17.9 11.0-11.0 12.1-12.1 4.7-4.7 15.9-15.9 8.6-8.6
Qwen Series Qwen2.5-7B-Instruct $42.6$ $58.2$ 35.3-35.3 21.7-21.7 50.7-50.7 22.0-22.0 35.7-35.7 28.0-28.0 37.7-37.7 7.0-7.0 39.8-39.8 19.7-19.7
Qwen2.5-14B-Instruct $48.4$ $70.7$ 36.1-36.1 22.9-22.9 37.3-37.3 35.3-35.3 33.0-33.0 43.4-43.4 14.2-14.2 10.9-10.9 30.2-30.2 28.1-28.1
Qwen2.5-32B-Instruct $59.9$ $79.4$ 25.4-25.4 17.3-17.3 23.5-23.5 18.5-18.5 23.6-23.6 24.1-24.1 19.7-19.7 5.4-5.4 23.1-23.1 16.3-16.3
Qwen2.5-72B-Instruct $58.1$ $82.8$ 24.0-24.0 12.3-12.3 15.4-15.4 6.9-6.9 24.9-24.9 16.3-16.3 25.2-25.2 6.5-6.5 22.4-22.4 10.5-10.5
Qwen2.5-Coder-32B-Instruct $63.1$ $84.8$ 24.4-24.4 14.6-14.6 28.2-28.2 19.0-19.0 21.7-21.7 19.9-19.9 14.7-14.7 9.3-9.3 22.3-22.3 15.7-15.7
Average 24.6-24.6 14.5-14.5 24.3-24.3 14.9-14.9 23.8-23.8 19.5-19.5 20.1-20.1 6.2-6.2 23.2-23.2 13.8-13.8
$\mathbf{-20.9}$ $\mathbf{-20.8}$ $\mathbf{-22.2}$ $\mathbf{-14.9}$ $\mathbf{-19.7}$
Table 15: Relative Pass@1 degradation of LLMs under all perturbations for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Absolute Version].
Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
@@
GPT-4 Omni GPT-4o-Mini $79.9$ $84.1$ $10.8\%$ $16.9\%$ $10.8\%$ $10.9\%$ $13.1\%$ $7.4\%$ $3.3\%$ $1.7\%$ $9.4\%$
GPT-4o $90.2$ $94.4$ $2.8\%$ $8.4\%$ $3.6\%$ $7.3\%$ $4.7\%$ $7.1\%$ $2.5\%$ $0.4\%$ $4.2\%$
Claude Series Claude-3.5-Haiku-20241022 $71.6$ $75.2$ $16.1\%$ $29.7\%$ $9.9\%$ $11.7\%$ $9.8\%$ $6.9\%$ $11.7\%$ $18.3\%$ $13.7\%$
Claude-3.5-Sonnet-20241022 $83.8$ $89.8$ $4.3\%$ $12.6\%$ $6.0\%$ $4.2\%$ $4.5\%$ $9.3\%$ $6.1\%$ $10.7\%$ $6.7\%$
Gemini Series Gemini-1.5-Flash $74.6$ $76.2$ $17.1\%$ $20.3\%$ $13.2\%$ $35.6\%$ $34.0\%$ $55.6\%$ $0\% 2.2$$2\%$ $21.0\%$
Gemini-2.0-Flash $87.2$ $92.3$ $5.0\%$ $8.1\%$ $6.0\%$ $6.8\%$ $7.6\%$ $24.9\%$ $2.3\%$ $1.6\%$ $7.1\%$
Gemini-1.5-Pro-002 $84.5$ $91.6$ $6.5\%$ $8.9\%$ $10.8\%$ $13.7\%$ $12.0\%$ $18.9\%$ $4.4\%$ $3.0\%$ $9.4\%$
DeepSeek DeepSeek-V3 $88.0$ $92.1$ $7.4\%$ $14.1\%$ $13.9\%$ $23.8\%$ $15.8\%$ $23.6\%$ $4.3\%$ $2.3\%$ $12.4\%$
LLaMA Series LLaMA-3.1-8B-Instruct $46.8$ $41.3$ $28.3\%$ $26.3\%$ $19.5\%$ $3.7\% 22$$2.2\%$ $20.2\%$ $5.6\%$ $14.6\%$ $19.8\%$
LLaMA-3.1-70B-Instruct $70.8$ $67.2$ $15.2\%$ $22.7\%$ $18.9\%$ $30.7\%$ $24.6\%$ $41.9\%$ $5.7\%$ $8.7\%$ $19.8\%$
LLaMA-3.1-405B-Instruct $78.9$ $77.7$ $9.8\%$ $19.9\%$ $7.3\%$ $15.3\%$ $10.1\%$ $24.7\%$ $4.0\%$ $3.4\% 11$$1.8\%$
LLaMA-3.3-70B-Instruct $76.1$ $78.1$ $8.2\%$ $15.2\%$ $5.4\%$ $11.5\%$ $5.7\%$ $19.8\%$ $1.8\%$ $9.6\%$ $8.6\%$
Qwen Series Qwen2.5-7B-Instruct $57.2$ $59.7$ $17.5\%$ $28.7\%$ $23.8\%$ $18.9\%$ $30.3\%$ $24.1\%$ $7.4\%$ $6.3\%$ $19.7\%$
Qwen2.5-14B-Instruct $67.2$ $76.4$ $20.3\%$ $27.3\%$ $36.1\%$ $34.2\%$ $43.5\%$ $43.2\%$ $11.7\%$ $9.6\%$ $28.1\%$
Qwen2.5-32B-Instruct $77.6$ $82.5$ $14.0\%$ $22.8\%$ $16.3\%$ $2.3\% 22$$2.4\%$ $27.1\%$ $6.6\%$ $3.3\%$ $16.3\%$
Qwen2.5-72B-Instruct $78.2$ $90.4$ $8.3\%$ $18.9\%$ $7.7\%$ $5.5\%$ $16.6\%$ $15.7\%$ $5.3\%$ $8.5\%$ $10.5\%$
Qwen2.5-Coder-32B-Instruct $84.1$ $85.8$ $10.1\%$ $22.1\%$ $16.8\%$ $22.6\%$ $14.0\%$ $29.7\%$ $8.0\%$ $11.4\%$ $15.7\%$
Average $11.9\%$ $19.0\%$ $13.3\%$ $17.6\%$ $17.1\%$ $23.5\%$ $5.5\%$ $7.3\%$
$\mathbf{-14.5}\%$ $\mathbf{-14.9}\%$ $\mathbf{-19.5}\%$ $\mathbf{-6.2}\%$ $\mathbf{-13.8}\%$
Table 16: Absolute Pass@1 degradation of LLMs under all perturbations for output prediction tasks across Crux and LCB in CoT inference ($N=1$). [To Relative Version].
Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
@@
GPT-4 Omni GPT-4o-Mini $79.9$ $84.1$ 8.6-8.6 14.2-14.2 8.6-8.6 9.2-9.2 10.5-10.5 6.3-6.3 2.6-2.6 1.5-1.5 7.7-7.7
GPT-4o $90.2$ $94.4$ 2.5-2.5 7.9-7.9 3.2-3.2 6.9-6.9 4.2-4.2 6.7-6.7 2.2-2.2 +0.4+0.4 3.9-3.9
Claude Series Claude-3.5-Haiku-20241022 $71.6$ $75.2$ 11.5-11.5 22.3-22.3 7.1-7.1 8.8-8.8 7.0-7.0 5.2-5.2 8.4-8.4 13.8-13.8 10.0-10.0
Claude-3.5-Sonnet-20241022 $83.8$ $89.8$ 3.6-3.6 11.3-11.3 5.0-5.0 3.8-3.8 3.8-3.8 8.4-8.4 5.1-5.1 9.6-9.6 5.8-5.8
Gemini Series Gemini-1.5-Flash $74.6$ $76.2$ 12.8-12.8 15.4-15.4 9.9-9.9 27.1-27.1 25.4-25.4 42.4-42.4 1.5-1.5 1.7-1.7 15.9-15.9
Gemini-2.0-Flash $87.2$ $92.3$ 4.4-4.4 7.5-7.5 5.2-5.2 6.3-6.3 6.6-6.6 23.0-23.0 2.0-2.0 1.5-1.5 6.4-6.4
Gemini-1.5-Pro-002 $84.5$ $91.6$ 5.5-5.5 8.1-8.1 9.1-9.1 12.5-12.5 10.1-10.1 17.3-17.3 3.8-3.8 2.7-2.7 8.3-8.3
DeepSeek DeepSeek-V3 $88.0$ $92.1$ 6.5-6.5 12.9-12.9 12.2-12.2 21.9-21.9 13.9-13.9 21.7-21.7 3.8-3.8 2.1-2.1 11.2-11.2
LLaMA Series LLaMA-3.1-8B-Instruct $46.8$ $41.3$ 13.2-13.2 10.9-10.9 9.1-9.1 9.8-9.8 10.4-10.4 8.4-8.4 2.6-2.6 6.1-6.1 8.8-8.8
LLaMA-3.1-70B-Instruct $70.8$ $67.2$ 10.8-10.8 15.2-15.2 13.4-13.4 20.7-20.7 17.4-17.4 28.2-28.2 4.0-4.0 5.8-5.8 13.7-13.7
LLaMA-3.1-405B-Instruct $78.9$ $77.7$ 7.8-7.8 15.4-15.4 5.8-5.8 11.9-11.9 8.0-8.0 19.2-19.2 3.1-3.1 10.4-10.4 9.2-9.2
LLaMA-3.3-70B-Instruct $76.1$ $78.1$ 6.2-6.2 11.9-11.9 4.1-4.1 9.0-9.0 4.4-4.4 15.4-15.4 1.4-1.4 7.5-7.5 6.6-6.6
Qwen Series Qwen2.5-7B-Instruct $57.2$ $59.7$ 10.0-10.0 17.1-17.1 13.6-13.6 11.3-11.3 17.4-17.4 14.4-14.4 4.2-4.2 3.8-3.8 11.4-11.4
Qwen2.5-14B-Instruct $67.2$ $76.4$ 13.6-13.6 20.9-20.9 24.2-24.2 26.1-26.1 29.2-29.2 33.0-33.0 7.9-7.9 7.3-7.3 19.9-19.9
Qwen2.5-32B-Instruct $77.6$ $82.5$ 10.9-10.9 18.8-18.8 12.6-12.6 18.4-18.4 17.4-17.4 22.3-22.3 5.1-5.1 2.7-2.7 13.0-13.0
Qwen2.5-72B-Instruct $78.2$ $90.4$ 6.5-6.5 17.1-17.1 6.0-6.0 5.0-5.0 13.0-13.0 14.2-14.2 4.1-4.1 7.7-7.7 8.8-8.8
Qwen2.5-Coder-32B-Instruct $84.1$ $85.8$ 8.5-8.5 19.0-19.0 14.1-14.1 19.4-19.4 11.8-11.8 25.5-25.5 6.8-6.8 9.8-9.8 13.3-13.3
Average 8.4-8.4 14.5-14.5 9.6-9.6 13.4-13.4 12.4-12.4 18.3-18.3 4.0-4.0 5.5-5.5
$\mathbf{-10.7}$ $\mathbf{-11.0}$ $\mathbf{-14.6}$ $\mathbf{-4.6}$ $\mathbf{-10.2}$
E.3 Additional Performance Degradation Tables for Input Prediction
In §3.5, we evaluate LLM robustness on input prediction using the Crux dataset. We focus on relative Pass@1 degradation (Table 5) to capture performance variation under perturbations. For completeness, we also report absolute degradation in Table 17.
Table 17: Absolute Pass@1 degradation for input prediction tasks across different reasoning modes on Crux. [To Relative Version].
Model Series Model Name VAN PSC-ALL MCC MPS MHC